# Modelling conductive thermal transport in three-dimensional fibrous media with fiber-to-fiber contacts

### v1

@@ II.2 Nodal analysis for the resolution of a steady-state conductive heat transport problem
In order to lower the uncertainty of the simulation results, the selected value of the solid thermal conductivity for a defined set of parameters is the mean of $k_{cond,solid}$ over 5 realizations; its associated 95% confidence interval is also computed and is represented under the form of error bars in all figures shown in this paper. This confidence level is a common choice in statistics: it has been used in particular in previous studies on similar fibrous systems, with $n\approx 10$ realizations [51, 56]. In this present work, errors are calculated using Student’s probability distribution, taking into account the fact that the standard deviations of samples of $n=5$ realizations underestimate the true Gaussian standard deviations. The relative error $\epsilon_{rel}$ on $k_{cond,solid}$ for each simulation point is obtained from the following expression: $\epsilon_{rel}=\frac{t_{n-1}(1-\alpha/2)s}{\overline{k_{cond,solid}}\sqrt{n}}$ where $\alpha$ is the significance level: here, $\alpha=0.05$ to achieve a 95% confidence interval. In this expression, $\overline{k_{cond,solid}}$ is the sample mean, $s$ is the unbiased estimator of the standard deviation given by $s=\sqrt{\frac{\sum_{i=1}^{n}(k_{cond,solid}^{i}-\overline{k_{cond,solid}})^{2}}{n-1}}$, and $t_{n-1}(1-\alpha/2)=2.78$ is the critical value of the t-distribution for a 95% confidence level. $n-1$, the number of realizations minus one, represents the number of degrees of freedom, i.e, the number of values that are allowed to vary when estimating the statistic.
@@
For each set of parameters, the sample volume $V=L^{3}$ is chosen so as to exceed the RVE, on top of the previous condition $L>l$. In this study, the RVE is defined as the volume for which the relative error on the solid thermal conductivity $\epsilon_{rel}$ introduced above is smaller than $5\%$ for $R_{k}=0$. This relative error criterion has been selected as similar approach was used by Altendorf et al. [51] for the selection of RVE in fibrous media; it also ensures a reasonable computation time. The procedure for selecting $L$ is illustrated in Fig. S2 (Supplementary Material) [50]: starting from a domain size slightly larger than the fiber length, it is gradually increased until the average of $k_{cond,solid}$ for $R_{k}=0$ on 5 realizations satisfies the error criteria $\epsilon_{rel}<5\%$, which also corresponds to achieving convergence of results. This ensures that the simulation results represent the whole network and are not dominated by the behavior of individual fibers. As a consequence of this procedure, the chosen value of $L$ depends on the input parameters; this is represented in Fig. S3 (Supplementary Material) [50]. To reach the RVE, a larger domain size is required when the volume fraction decreases, or when fiber dimensions increase, all other parameters being fixed, which can be attributed to the decrease of the total number of fibers. In contrast, the average fiber angle appears to have only a minor impact on the selection of $L$.
@@
Numerical simulations are performed to independently study the influence of geometric fiber parameters on the resulting thermal properties of the network. In particular, the joint effects of fiber volume fraction $V_{f}$, fiber length $l$, fiber diameter $d$ and fiber orientation $\bar{\theta}$ with contact resistance $R_{k}$ are investigated. For each set of parameters and on top of the conditions on $V$ and on $L$ described previously, as some fibers are deleted in the generation process, the orientation and length distributions are compared between the initial and final states, to ensure they are not substantially modified. For all simulations in this study, the size of the calculation domain ranges from $L=1.25$ mm to $L=6$ mm depending on the four parameters, the contact density from 40 to 20000 mm-3, the average number of contacts per fiber from 2 to 20, and the fiber density from 25 to 4000 mm-3.
@@ III.1.1 Effect of thermal contact resistance at different fiber volume fractions
The variation of solid conductivity with fiber volume fraction $V_{f}$, fixing fiber orientation, length and diameter, is illustrated in Fig. 4
Figure 4: (a) Solid conductivity as a function of fiber volume fraction for different thermal contact resistances. The dashed lines indicate the power laws best fits to the simulation data for $0.03\geq V_{f}\geq 0.1$, with exponents varying ranging from 1.25 to 2.23 as the contact resistance increases from 00 to $10^{9}$ KW-1. The constant input parameters are $l=1$ mm, $d=10~{}\mu$m and $\bar{\theta}=57.3^{\circ}$. (b) Plot of the corresponding contact densities, as a function of fiber volume fraction. The dashed line indicates the square power law best fit to the simulation data for $0.03\geq V_{f}\geq 0.1$.
(a) for different contact resistances. Volume fractions between 0.0075 and 0.1 are studied, corresponding to glasswool mat densities ranging from 17 kg.m-3 to 221 kg.m-3. In all cases, the solid conductivity increases with density in a non-linear way; in particular, it shows power-law behavior at high volume fractions ($V_{f}\geq 0.03$) indicated by the dashed line best-fits. The graph also shows that the increase in solid conductivity with volume fraction becomes more pronounced as the contact resistance increases, i.e., the exponent in the power law increases, from 1.25 for $R_{k}=0$ to 2.23 for $10^{9}$ KW-1. This generalizes previous results reported in Refs. [38] and [40] for 2D networks to the case of 3D networks. These studies have also distinguished two regimes depending on volume fraction: the low-density regime, close to the percolation threshold, exhibits a different behavior from that observed at higher densities. This is presumably due to a difference in the degree of connectivity of the networks, i.e., in the proportion of fibers and contacts conveying heat. Indeed, the contact density increases when the volume fraction increases, as shown in Fig. 4(b). For high densities, it is expected to follow a square power law [38], which is verified in our simulation results for $V_{f}\geq 0.03$. This value of volume fraction provides a first estimate of the high-density threshold for the chosen simulation parameters, consistent with the validity range of the power laws introduced in Fig. 4(a). According to this estimation, networks with $V_{f}=0.01$ should correspond to a low-density regime in which the connectivity is low: to simplify the study of the impact of the other fiber parameters, this value of volume fraction will be chosen in the following sections.
@@ III.1.2 Effect of thermal contact resistance at different fiber orientations
Fig. 5
Figure 5: (a) Solid conductivity as a function of the mean angle to heat flux direction for different thermal contact resistances. The anisotropy parameter $\beta$ ranges from 0.1 to 30, corresponding to mean angles $\bar{\theta}$ between 31.8 and 85.5: only data obtained for $\beta>1$, i.e., for networks in which fibers are preferentially oriented in the cross-plane direction, are representative of insulation materials such as glasswool. The constant input parameters are $l=1$ mm, $d=10~{}\mu$m and $V_{f}=0.01$. (b) Representation of simplified networks illustrating a qualitative argument for the dependence of the conductivity on the mean angle to heat flux direction: when it increases, the number of conduction paths decreases. (c) Plot of contact densities, corresponding to the simulations of (a), as a function of the mean angle.
(a) illustrates the variation of solid conductivity as a function of the mean angle between the fibers and the heat flux direction $\bar{\theta}$, for different values of contact resistance and for fixed fiber length, diameter and volume fraction. The mean angle $\bar{\theta}=57.3^{\circ}$ corresponds to a uniform distribution in spherical coordinates, i.e., to $\beta=1$; then $\beta$ is increased up to 30, corresponding to a mean angle of 85.5, so as to cover the estimated range for $\bar{\theta}$ in glasswool, between 75 and 85 [53]. For the sake of completeness, we have also included in this figure the simulation results obtained for $\beta<1$. As $\beta$ decreases below 1 and approaches 0, fibers become increasingly aligned approximately parallel to the $z$-axis, as it can be seen in Figs. 1 and 3. Note that this case is not representative of insulation materials, in which fibers tend to form significant angles to the heat flux direction. The graph shows that the solid conductivity decreases as the mean angle to the heat flux direction increases whether contact resistances are integrated or not. This indicates that the impact of contacts on solid conductivity does not change with the orientation of fibers, including when they tend to align with the heat propagation axis.
Note that at sufficiently low densities, the interconnected conduction paths are few in number and can be neglected as a first approximation. In this case, a simplified expression for the total thermal resistance can be built based on simplified networks such as those represented in Fig. 5(b), accounting for the contributions of several conduction paths in parallel, each consisting of several fibers in series. The total resistance $R_{t}$ of such networks is the ratio of two parameters: the average resistance of the paths $R_{p}$ and the number of conductive paths in parallel $n_{p}$ [57, 58]. This average path resistance $R_{p}$ is the result of two factors that add up as a series: the total resistance of contacts, and the total internal resistance of the path. Therefore,
@@
where $n_{c,p}$ is the average number of contacts per path, $n_{c,tot}$ the total number of contacts, $l_{p}$ the average length of a path and $l_{tot}$ the total length of the paths. In these considerations, the unconnected fiber ends are neglected: therefore, it is assumed that $l_{tot}$ is the total length of fibers in the networks. As shown in Fig. 5(b), more fibers are needed to connect one heat bath to the other through one path when $\bar{\theta}$ increases. Therefore, the created paths are overall less direct and longer, leading to a decrease in the total number of paths $n_{p}$. As the volume fraction and the fiber dimensions are fixed in the simulations, the total fiber length is constant: Eq. (2) confirms that $R_{t}$ increases with $\bar{\theta}$ when there is no contact resistance, leading to a decrease in the conductivity. Moreover, the parameter governing the relative impact of contacts with respect to that of internal resistance is the total number of contacts $n_{c,tot}$. Fig. 5(c) shows that the volume density of contacts remains around 200 mm-3 regardless of fiber orientation for the simulation parameters chosen in this study: this supports our observation that the effects of contacts and of fiber mean angle on the network are not correlated in the studied range, and particularly in that representative of glasswool.
@@ III.1.3 Effect of thermal contact resistance at different fiber lengths
Fig. 6
Figure 6: (a) Solid conductivity as a function of fiber length for different thermal contact resistances. The constant input parameters are $\bar{\theta}=57.3^{\circ}$, $d=10~{}\mu$m and $V_{f}=0.01$. (b) Representation of simplified networks illustrating a qualitative argument for the dependence of the conductivity on fiber length: when it increases, the number of conduction paths increases and the number of contacts per path decreases [59]. (c) Plot of contact densities, corresponding to the simulations of (a), as a function of fiber length.
(a) displays a plot of solid conductivity versus fiber length $l$ for different contact resistances, fixing the other fiber parameters. Fiber length ranges from 0.75 mm to 5 mm in the simulations, remaining within the same order of magnitude of 1 mm, which is that representative of glasswool [53]. The graph shows that when no contact resistance is considered, the solid conductivity increases when fiber length increases. Including the impact of junctions between fibers, our results show that an increasing contact resistance induces a decrease in the solid conductivity compared with its value at $R_{k}=0$, and that the impact of fiber length is amplified when the contact resistance increases. This means that contact resistance has a lower relative impact for longer fiber assemblies. This behavior is in agreement with the results obtained in Ref. [40] for electrically conducting 2D nanowire networks.
The density and diameter being fixed, the total length is constant and the number of fibers needed to create a path decreases when $l$ increases. Conduction paths are becoming less tortuous and shorter: as a consequence, more conduction paths in parallel can convey heat. According to Eq. (2), this explains the observed increase in the solid conductivity when there is no contact resistance; Fig. 6(b) illustrates this situation. Eq. (2) also shows that when it comes to the impact of contacts in simplified networks, the total number of contacts is the parameter that determines its relative importance with respect to the impact of internal resistance, similar to the case of fiber orientation variation. The evolution of contact density versus fiber length is presented in Fig. 6(c). For $l<2.5$ mm, contact density decreases with fiber length. Then, it seems to reach a threshold - the onset of decline observed at $l>4.5$ mm remains to be confirmed by further simulations. This behavior is in accordance with previously established theoretical models that estimate the number of contacts in a network of lines [60]. Overall, we find that contact density tends to decrease over the studied range. According to Eq. (2) and as represented in the examples of Fig. 6(b), this should lead to a reduction of the resistive contribution of contacts in the network as there are fewer contacts per path, which is consistent with the correlation observed in Fig. 6(a) between impact of contacts and impact of fiber length.
@@ III.1.4 Effect of thermal contact resistance at different fiber diameters
Figure 7: (a) Solid conductivity as a function of fiber diameter for different thermal contact resistances. The constant input parameters are $\bar{\theta}=57.3^{\circ}$, $l=1$ mm and $V_{f}=0.01$. (b) Representation of simplified networks illustrating a qualitative argument for the dependence of conductivity on fiber diameter: when it increases, the number of conduction paths in parallel decreases. Assuming a constant number of contacts per path, the number of contacts resistances $R_{k}$ in parallel decreases, leading to an increase in the total resistive contribution of junctions, which goes from $R_{k}/4$ to $R_{k}$ in this example [59]. (c) Plot of contact densities, corresponding to the simulations of (a), as a function of fiber diameter.
@@
(a). The diameter ranges from 2 $\mu$m to 20 $\mu$m as so as to cover the sizes of the glass fibers that can be found in real glasswool mats [53]. The graph shows that with no contact resistance, the solid conductivity decreases when fiber diameter increases. Also, similar to the case of fiber length, the impact of fiber diameter is amplified when the contact resistance increases.
@@
Our numerical results indicate that among the four ranges of variation for fiber parameters studied individually, the one chosen for fiber diameter leads to the greatest variation in the total number of fibers. This means that changing the diameter between 2 $\mu$m and 20 $\mu$m greatly affects the connectivity of the network. As a consequence, the assumption of low-density in the sense of no interconnection between conduction paths cannot be made here over the whole diameter range, even though the volume fraction is fixed at 0.01. Networks with a higher diameter are composed of less fibers, and therefore exhibit less connectivity, leading to a less efficient heat flux distribution which explains the observed decrease in the solid conductivity. This effect is expected to be important as the fixed volume fraction is low [61]. Moreover, once again Eqs. (1) or (2) give insights about the evolution of the relative impact of contacts with respect to that of the internal resistance with fiber diameter, which is illustrated by simplified networks in Fig. 7(b). The total contact density decreases as the fiber diameter increases – this is verified by simulation results shown in Fig. 7(c), and the number of conduction paths in parallel decreases as well. From Eq. (1), assuming constant $l_{p}$ and $n_{c,p}$, an increase in fiber diameter increases the ratio of the resistive contribution of contacts over that of paths due to its dependence on $d^{2}$. This leads to an amplification of the impact of the remaining contacts resistances, even though there are fewer of them. These considerations, although there are limited to a simplified case with no interconnection, are consistent with the correlation between impact of contacts and impact of fiber diameter observed in Fig. 7(a).
@@
Note that while in this study, contact resistance and fiber diameter are treated as independent, it is likely that a correlation exists between these two parameters [59]. Indeed, glass fibers are expected to deform at a junction, according to the Hertz contact law. Therefore, the contact area between fibers with a larger diameter would increase, leading to a lower associated contact resistance. In the configuration studied here, that can be described as a perfectly rigid assumption, one can notice that the effects of fiber length and diameter seem to have inverse effects, indicating a potential dependence of the network properties on the aspect ratio $a$. In particular, the influence of aspect ratio on the solid conductivity is amplified in the case of resistive contacts, which was also reported in Ref. [42].

### v2

@@ II.2 Nodal analysis for the resolution of a steady-state conductive heat transport problem
In order to lower the uncertainty of the simulation results, the selected value of the solid thermal conductivity for a defined set of parameters is the mean of $k_{cond,solid}$ over five realizations; its associated 95% confidence interval is also computed and is represented under the form of error bars in all figures shown in this paper. This confidence level is a common choice in statistics: it has been used in particular in previous studies on similar fibrous systems, with $n\approx 10$ realizations [51, 56]. In this work, errors are calculated using the Student’s probability distribution, taking into account the fact that the standard deviations of samples of $n=5$ realizations underestimate the true Gaussian standard deviations. The relative error $\epsilon_{rel}$ on $k_{cond,solid}$ for each simulation point is obtained from the expression $\epsilon_{rel}=\frac{t_{n-1}(1-\alpha/2)s}{\overline{k_{cond,solid}}\sqrt{n}}$ where $\alpha$ is the significance level; here $\alpha=0.05$ to achieve a 95% confidence interval. In this expression, $\overline{k_{cond,solid}}$ is the sample mean, $s$ is the unbiased estimator of the standard deviation given by $s=\sqrt{\frac{\sum_{i=1}^{n}(k_{cond,solid}^{i}-\overline{k_{cond,solid}})^{2}}{n-1}}$, and $t_{n-1}(1-\alpha/2)=2.78$ is the critical value of the $t$ distribution for a 95% confidence level. Here $n-1$, the number of realizations minus one, represents the number of degrees of freedom, i.e, the number of values that are allowed to vary when estimating the statistic.
@@
For each set of parameters, the sample volume $V=L^{3}$ is chosen so as to exceed the RVE, on top of the previous condition $L>l$. In this study, the RVE is defined as the volume for which the relative error on the solid thermal conductivity $\epsilon_{rel}$ introduced above is smaller than $5\%$ for $R_{k}=0$. This relative error criterion has been selected as a similar approach was used by Altendorf et al. [51] for the selection of RVE in fibrous media; it also ensures a reasonable computation time. The procedure for selecting $L$ is illustrated in Fig. S2 within the Supplemental Material [50]: starting from a domain size slightly larger than the fiber length, it is gradually increased until the average of $k_{cond,solid}$ for $R_{k}=0$ on five realizations satisfies the error criteria $\epsilon_{rel}<5\%$, which also corresponds to achieving convergence of the results. This ensures that the simulation results represent the whole network and are not dominated by the behavior of individual fibers. As a consequence of this procedure, the chosen value of $L$ depends on the input parameters; this is represented in Fig. S3 within the Supplemental Material [50]. To reach the RVE, a larger domain size is required when the volume fraction decreases, or when fiber dimensions increase, all other parameters being fixed, which can be attributed to the decrease of the total number of fibers. In contrast, the average fiber angle appears to have only a minor impact on the selection of $L$.
@@
Numerical simulations are performed to independently study the influence of geometric fiber parameters on the resulting thermal properties of the network. In particular, the joint effects of the fiber volume fraction $V_{f}$, fiber length $l$, fiber diameter $d$, and fiber orientation $\bar{\theta}$ with contact resistance $R_{k}$ are investigated. For each set of parameters and on top of the conditions on $V$ and on $L$ described previously, as some fibers are deleted in the generation process, the orientation and length distributions are compared between the initial and final states, to ensure that they are not substantially modified. For all simulations in this study, the size of the calculation domain ranges from $L=1.25$ mm to $L=6$ mm depending on the four parameters, the contact density ranges from 40 to 20000 mm-3, the average number of contacts per fiber ranges from 2 to 20, and the fiber density ranges from 25 to 4000 mm-3.
@@ III.1.1 Effect of thermal contact resistance at different fiber volume fractions
The variation of the solid conductivity with the fiber volume fraction $V_{f}$, fixing the fiber orientation, length, and diameter, is illustrated in Fig. 4
Figure 4: (a) Solid conductivity as a function of the fiber volume fraction for different thermal contact resistances. The dashed lines indicate the power-law best fits to the simulation data for $0.03\geq V_{f}\geq 0.1$, with exponents ranging from 1.25 to 2.23 as the contact resistance increases from 00 to $10^{9}$ K W-1. The constant input parameters are $l=1$ mm, $d=10~{}\mu$m, and $\bar{\theta}=57.3^{\circ}$. (b) Plot of the corresponding contact densities, as a function of fiber volume fraction. The dashed line indicates the square power-law best fit to the simulation data for $0.03\geq V_{f}\geq 0.1$.
(a) for different contact resistances. Volume fractions between 0.0075 and 0.1 are studied, corresponding to glasswool mat densities ranging from 17 to 221 kg m-3. In all cases, the solid conductivity increases with density in a nonlinear way; in particular, it shows power-law behavior at high volume fractions ($V_{f}\geq 0.03$) indicated by the dashed line best fits. The graph also shows that the increase in the solid conductivity with the volume fraction becomes more pronounced as the contact resistance increases, i.e., the exponent in the power law increases, from 1.25 for $R_{k}=0$ to 2.23 for $10^{9}$ K W-1. This generalizes previous results reported in Refs. [38] and [40] for 2D networks to the case of 3D networks. These studies have also distinguished two regimes depending on volume fraction: the low-density regime, close to the percolation threshold, exhibits a different behavior from that observed at higher densities. This is presumably due to a difference in the degree of connectivity of the networks, i.e., in the proportion of fibers and contacts conveying heat. Indeed, the contact density increases when the volume fraction increases, as shown in Fig. 4(b). For high densities, it is expected to follow a square power law [38], which is verified in our simulation results for $V_{f}\geq 0.03$. This value of the volume fraction provides a first estimate of the high-density threshold for the chosen simulation parameters, consistent with the validity range of the power laws introduced in Fig. 4(a). According to this estimation, networks with $V_{f}=0.01$ should correspond to a low-density regime in which the connectivity is low: to simplify the study of the impact of the other fiber parameters, this value of the volume fraction will be chosen in the following sections.
@@ III.1.2 Effect of thermal contact resistance at different fiber orientations
Figure 5
Figure 5: (a) Solid conductivity as a function of the mean angle to heat flux direction for different thermal contact resistances. The anisotropy parameter $\beta$ ranges from 0.1 to 30, corresponding to mean angles $\bar{\theta}$ between 31.8 and 85.5: only data obtained for $\beta>1$, i.e., for networks in which fibers are preferentially oriented in the cross-plane direction, are representative of insulation materials such as glasswool. The constant input parameters are $l=1$ mm, $d=10~{}\mu$m, and $V_{f}=0.01$. (b) Representation of simplified networks illustrating a qualitative argument for the dependence of the conductivity on the mean angle to heat flux direction: when it increases, the number of conduction paths decreases. (c) Plot of contact densities, corresponding to the simulations of (a), as a function of the mean angle.
(a) illustrates the variation of the solid conductivity as a function of the mean angle between the fibers and the heat flux direction $\bar{\theta}$, for different values of the contact resistance and for fixed fiber length, diameter, and volume fraction. The mean angle $\bar{\theta}=57.3^{\circ}$ corresponds to a uniform distribution in spherical coordinates, i.e., to $\beta=1$; then $\beta$ is increased up to 30, corresponding to a mean angle of 85.5, so as to cover the estimated range for $\bar{\theta}$ in glasswool, between 75 and 85 [53]. For the sake of completeness, we have also included in this figure the simulation results obtained for $\beta<1$. As $\beta$ decreases below 1 and approaches 0, fibers become increasingly aligned approximately parallel to the $z$ axis, as can be seen in Figs. 1 and 3. Note that this case is not representative of insulation materials, in which fibers tend to form significant angles to the heat flux direction. The graph shows that the solid conductivity decreases as the mean angle to the heat flux direction increases whether contact resistances are integrated or not. This indicates that the impact of contacts on the solid conductivity does not change with the orientation of fibers, including when they tend to align with the heat propagation axis.
Note that at sufficiently low densities, the interconnected conduction paths are few in number and can be neglected as a first approximation. In this case, a simplified expression for the total thermal resistance can be built based on simplified networks such as those represented in Fig. 5(b), accounting for the contributions of several conduction paths in parallel, each consisting of several fibers in series. The total resistance $R_{t}$ of such networks is the ratio of two parameters: the average resistance of the paths, $R_{p}$, and the number of conductive paths in parallel, $n_{p}$ [57, 58]. This average path resistance $R_{p}$ is the result of two factors that add up as a series: the total resistance of contacts, and the total internal resistance of the path. Therefore,
@@
where $n_{c,p}$ is the average number of contacts per path, $n_{c,tot}$ is the total number of contacts, $l_{p}$ is the average length of a path and $l_{tot}$ is the total length of the paths. In these considerations, the unconnected fiber ends are neglected: therefore, it is assumed that $l_{tot}$ is the total length of fibers in the networks. As shown in Fig. 5(b), more fibers are needed to connect one heat bath to the other through one path when $\bar{\theta}$ increases. Therefore, the created paths are overall less direct and longer, leading to a decrease in the total number of paths $n_{p}$. As the volume fraction and the fiber dimensions are fixed in the simulations, the total fiber length is constant: Eq. (2) confirms that $R_{t}$ increases with $\bar{\theta}$ when there is no contact resistance, leading to a decrease in the conductivity. Moreover, the parameter governing the relative impact of contacts with respect to that of internal resistance is the total number of contacts $n_{c,tot}$. Figure 5(c) shows that the volume density of contacts remains around 200 mm-3 regardless of fiber orientation for the simulation parameters chosen in this study: this supports our observation that the effects of contacts and of the fiber mean angle on the network are not correlated in the studied range, and particularly in that representative of glasswool.
@@ III.1.3 Effect of thermal contact resistance at different fiber lengths
Figure 6
Figure 6: (a) Solid conductivity as a function of fiber length for different thermal contact resistances. The constant input parameters are $\bar{\theta}=57.3^{\circ}$, $d=10~{}\mu$m, and $V_{f}=0.01$. (b) Representation of simplified networks illustrating a qualitative argument for the dependence of the conductivity on fiber length: when it increases, the number of conduction paths increases and the number of contacts per path decreases [59]. (c) Plot of contact densities, corresponding to the simulations of (a), as a function of fiber length.
(a) displays a plot of the solid conductivity versus the fiber length $l$ for different contact resistances, fixing the other fiber parameters. The fiber length ranges from 0.75 to 5 mm in the simulations, remaining within the same order of magnitude of 1 mm, which is representative of glasswool [53]. The graph shows that, when no contact resistance is considered, the solid conductivity increases when the fiber length increases. Including the impact of junctions between fibers, our results show that an increasing contact resistance induces a decrease in the solid conductivity compared with its value at $R_{k}=0$, and that the impact of fiber length is amplified when the contact resistance increases. This means that contact resistance has a lower relative impact for longer fiber assemblies. This behavior is in agreement with the results obtained in Ref. [40] for electrically conducting 2D nanowire networks.
The density and diameter being fixed, the total length is constant and the number of fibers needed to create a path decreases when $l$ increases. Conduction paths are becoming less tortuous and shorter: as a consequence, more conduction paths in parallel can convey heat. According to Eq. (2), this explains the observed increase in the solid conductivity when there is no contact resistance; Fig. 6(b) illustrates this situation. Equation (2) also shows that when it comes to the impact of contacts in simplified networks, the total number of contacts is the parameter that determines its relative importance with respect to the impact of internal resistance, similar to the case of fiber orientation variation. The evolution of contact density versus fiber length is presented in Fig. 6(c). For $l<2.5$ mm, contact density decreases with fiber length. Then, it seems to reach a threshold - the onset of decline observed at $l>4.5$ mm remains to be confirmed by further simulations. This behavior is in accordance with previously established theoretical models that estimate the number of contacts in a network of lines [60]. Overall, we find that contact density tends to decrease over the studied range. According to Eq. (2) and as represented in the examples of Fig. 6(b), this should lead to a reduction of the resistive contribution of contacts in the network as there are fewer contacts per path, which is consistent with the correlation observed in Fig. 6(a) between the impact of contacts and the impact of fiber length.
@@ III.1.4 Effect of thermal contact resistance at different fiber diameters
Figure 7: (a) Solid conductivity as a function of fiber diameter for different thermal contact resistances. The constant input parameters are $\bar{\theta}=57.3^{\circ}$, $l=1$ mm, and $V_{f}=0.01$. (b) Representation of simplified networks illustrating a qualitative argument for the dependence of conductivity on fiber diameter: when it increases, the number of conduction paths in parallel decreases. Assuming a constant number of contacts per path, the number of contact resistances $R_{k}$ in parallel decreases, leading to an increase in the total resistive contribution of junctions, which goes from $R_{k}/4$ to $R_{k}$ in this example [59]. (c) Plot of contact densities, corresponding to the simulations of (a), as a function of fiber diameter.
@@
(a). The diameter ranges from 2 to 20 $\mu$m so as to cover the sizes of the glass fibers that can be found in real glasswool mats [53]. The graph shows that, with no contact resistance, the solid conductivity decreases when the fiber diameter increases. Also, similar to the case of fiber length, the impact of fiber diameter is amplified when the contact resistance increases.
@@
Our numerical results indicate that among the four ranges of variation for fiber parameters studied individually, the one chosen for the fiber diameter leads to the greatest variation in the total number of fibers. This means that changing the diameter between 2 and 20 $\mu$m greatly affects the connectivity of the network. As a consequence, the assumption of low density in the sense of no interconnection between conduction paths cannot be made here over the whole diameter range, even though the volume fraction is fixed at 0.01. Networks with a higher diameter are composed of less fibers, and therefore exhibit less connectivity, leading to a less efficient heat flux distribution which explains the observed decrease in the solid conductivity. This effect is expected to be important as the fixed volume fraction is low [61]. Moreover, once again Eq. (1) or (2) gives insights into the evolution of the relative impact of contacts with respect to that of the internal resistance with fiber diameter, which is illustrated by simplified networks in Fig. 7(b). The total contact density decreases as the fiber diameter increases – this is verified by simulation results shown in Fig. 7(c), and the number of conduction paths in parallel decreases as well. From Eq. (1), assuming constant $l_{p}$ and $n_{c,p}$, an increase in the fiber diameter increases the ratio of the resistive contribution of contacts over that of paths due to its dependence on $d^{2}$. This leads to an amplification of the impact of the remaining contact resistances, even though there are fewer of them. These considerations, although they are limited to a simplified case with no interconnection, are consistent with the correlation between the impact of contacts and impact of fiber diameter observed in Fig. 7(a).
@@
Note that while in this study contact resistance and fiber diameter are treated as independent, it is likely that a correlation exists between these two parameters [59]. Indeed, glass fibers are expected to deform at a junction, according to the Hertz contact law. Therefore, the contact area between fibers with a larger diameter would increase, leading to a lower associated contact resistance. In the configuration studied here, which can be described as a perfectly rigid assumption, one can note that the effects of fiber length and diameter seem to have inverse effects, indicating a potential dependence of the network properties on the aspect ratio $a$. In particular, the influence of the aspect ratio on the solid conductivity is amplified in the case of resistive contacts, which was also reported in Ref. [42].